OpenStudy (anonymous):

Evaluate the following. Write your answer in scientific notation. (3 x 107)(6.1 x 103)

Parth (parthkohli):

=> (321)(628.3) => 201684.3 => 2.016843 * 10^5

OpenStudy (unklerhaukus):

\[3\times 10^7 \times 6.1 \times 10^3=18.3\times 10^{10}=1.83\times10^{11}\]